A soup can is made up of a side, a top, and a bottom. If the diameter and height of the can are equal and the volume is 128pi units^3, what is the total surface area of the can in square units?
Answer by Cromlix(4381)

Diameter = x
Radius = x/2
Height = x
Volume = Pi*(x/2)^2*x
128Pi = Pi*x/4 * x
128 = x^3/4
x^3 = 4 * 128
x^3 = 512
x = cube root of 512
x = 8 units
Height = 8 units
Radius = 4 units
Surface area of totally enclosed can =
2*Pi*r^2 + Pi*r*h
2*Pi*4^2 + Pi*4*8
=201.1 units^2
Hope this helps.
;-)